OpenStudy (anonymous):

Simplify the square root of 5 divided by 6. a. 6* sq rt of 5 divided by 5 b. 5* sq rt of 6 divided by 6 c. sq rt of 30 divided by 5 d. sq rt of 30 divided by 6

OpenStudy (lgbasallote):

\[\huge \sqrt{\frac56}?\]

OpenStudy (anonymous):

yes. how do you do the sq rt sign ? lol

OpenStudy (anonymous):

Do you know how to rationalize the denominator??

OpenStudy (lgbasallote):

\[\huge \sqrt{\frac56} \implies \frac{\sqrt 5}{\sqrt 6} \times \frac{\sqrt 6}{\sqrt 6} \implies \frac{\sqrt{5 \times 6}}{6}\]

OpenStudy (anonymous):

how do you do the sq rt symbol ? @lgbasallote

OpenStudy (lgbasallote):

you see the "Equation" button below the reply box? click it...you'll see a tab that has a \(\sqrt{}\) symbol

OpenStudy (anonymous):

well thank ya (:

OpenStudy (lgbasallote):

\[\Huge \color{maroon}{\mathtt{\text{<tips hat>}}}\]

OpenStudy (anonymous):

@lgbasallote : wait i cant do the symbol when im asking a question because it doesnt have an equation thing lol. so what do i do ?

OpenStudy (lgbasallote):

do exactly as i say type in \[\huge \text{\[\sqrt{radicand}\]}\] if you want an index.. \[\huge \text{\[\sqrt[index]{radicand}\]}\]

OpenStudy (anonymous):

whats an index ? can you show me and example of what both would look like ?please.

OpenStudy (lgbasallote):

\[\huge \text{\sqrt{radicand}} \implies \sqrt{radicand}\] \[\huge \text{\sqrt[index]{radicand}} \implies \sqrt[index]{radicand}\]

OpenStudy (lgbasallote):

\[\huge \text{\sqrt{m}} \implies \sqrt m\] \[\huge \text{\sqrt[n]{m}} \implies \sqrt[n]{m}\] got it?

OpenStudy (anonymous):

ok cool. yes!

OpenStudy (lgbasallote):

glad to hear that :)